# Numerical tiling-based simulations of decoherence in multifield models of inflation

### v2

@@ IV Discussion
In this project, we present a numerical implementation that incorporates the simplest non-unitary decoherence corrections into the dynamics of inflationary fluctuations at the level of two-point correlators. These corrections enable controlled manipulations of the initial Gaussian states in a way analogous to a pumping protocol in quantum optics. The method accommodates arbitrary arrangements of decoherence events and accounts for nonlinear background field evolution without relying on the slow-roll approximations. Furthermore, this approach has been extended to evaluate systems with multiple fields that undergo coupling through the time-evolving background fields. The numerical implementation employs the Wigner representation of Gaussian states to perform dynamical coloring transformations – which, as demonstrated in [28], act as a fast-slow scale separation scheme that efficiently suppresses the highest oscillation scales of a problem.

### v3

@@ IV Discussion
In this project, we present a numerical implementation that incorporates the leading non-unitary decoherence corrections into the dynamics of inflationary fluctuations at the level of two-point correlators. It is important to emphasize that the results presented here are intended to illustrate the diagnostic power of the framework, and should not be interpreted as predictions of any specific inflationary model.
These corrections enable controlled manipulations of the initial Gaussian states in a way analogous to a pumping protocol in quantum optics. The method accommodates arbitrary arrangements of decoherence events and accounts for nonlinear background field evolution without relying on the slow-roll approximations. Furthermore, this approach has been extended to evaluate systems with multiple fields that undergo coupling through the time-evolving background fields. The numerical implementation employs the Wigner representation of Gaussian states to perform dynamical coloring transformations – which, as demonstrated in [28], act as a fast-slow scale separation scheme that efficiently suppresses the highest oscillation scales of a problem.